OpenStudy (anonymous):

What is the radius of a circle with a circumference of twenty-six point six nine meters? Use three point one four for pi and show your work.

OpenStudy (anonymous):

c = 2 pi r 26.9 = 2(3.14) r 26.9 = 2.68 [I think re check it] r 26.9 / 2.68 = r :)

OpenStudy (anonymous):

c = 2 pi r 26.9 = 2(3.14) r 26.9 = 6.82 r 26.9 / 6.82 = r

OpenStudy (empty):

oh almost theres like two sixes there 26.69